OpenStudy (anonymous):

What is the rule for the sequence 5, 13, 21, 29, …? Add 8 to the previous number Multiply the previous number by 2 Add 5 to the previous number Multiply the previous number by 3

OpenStudy (anonymous):

add 8 to the previous number 5+8=13 13+8=21 21+8=29